# Beyond Semantics: Rediscovering Spatial Awareness in Vision-Language Models

### v1

@@ Abstract
Vision-Language Models (VLMs) excel at identifying and describing objects but struggle with spatial reasoning such as accurately understanding the relative positions of objects. Inspired by the dual-pathway (ventral-dorsal) model of human vision, we investigate why VLMs fail spatial tasks despite strong object recognition capabilities. Our interpretability-driven analysis reveals a critical underlying cause: vision embeddings in VLMs are treated primarily as semantic “bag-of-tokens,” overshadowing subtle yet crucial positional cues due to their disproportionately large embedding norms. We validate this insight through extensive diagnostic experiments, demonstrating minimal performance impact when token orders or fine-grained spatial details are removed. Guided by these findings, we propose simple, interpretable interventions, including normalizing vision embedding norms and extracting mid-layer spatially rich features, to restore spatial awareness. Empirical results on both our synthetic data and standard benchmarks demonstrate improved spatial reasoning capabilities, highlighting the value of interpretability-informed design choices. Our study not only uncovers fundamental limitations in current VLM architectures but also provides actionable insights for enhancing structured perception of visual scenes. Website is at: user074.github.io/respatialaware/
Figure 1: Permutation Test: Original (top) vs. randomly permuted vision tokens (bottom). Despite losing spatial ordering, the LLM accurately responds to the prompt “Describe the image,” demonstrating strong robustness and a notable “bag-of-tokens” tendency. Token embeddings are visualized using cosine similarity relative to a reference token.
@@ 1 Introduction
Humans process vision through two distinct pathways: the ventral stream, which identifies objects, and the dorsal stream, which encodes spatial relationships [10]. The dorsal stream, located in the parietal lobe, is essential for processing spatial relationships, allowing us to understand where objects are and how they relate to each other. Interestingly, patients with damage to the parietal lobe exhibit asymmetric visual processing: they retain the ability to recognize and describe objects but lose spatial reasoning capabilities [10].
@@
Recent observations suggest that Vision-Language Models (VLMs) exhibit a similar asymmetry. Despite remarkable performance in object-centric (ventral) tasks such as object recognition and image captioning[11, 2, 12], VLMs systematically struggle with even fundamental spatial reasoning queries (e.g., distinguishing “left” from “right”) [19, 33]. For example, when asked, “Is the apple to the left of the cup?”, current state-of-the-art VLMs can accurately identify both objects, yet frequently fail to correctly infer their spatial relationship. This leads to a critical interpretability question: Why do VLMs systematically fail at spatial reasoning even though they accurately recognize objects?
One possible reason is the inherent treatment of visual inputs in VLMs as a “bag-of-tokens”, analogous to the “bag-of-words” representation in NLP, where token order is disregarded [13, 38]. In this kind of representation, vision embeddings lose intrinsic spatial structure, significantly limiting relational awareness. However, the underlying interpretability mechanisms behind this phenomenon remain poorly understood.
We propose that adopting the ventral-dorsal analogy provides a useful interpretability heuristic to analyze how VLMs process what versus where information. Specifically, we hypothesize that spatial reasoning failures are exacerbated not merely by insufficient positional encoding but crucially by embedding norm suppression—the phenomenon where large embedding magnitudes overwhelm positional signals in attention mechanisms.
To rigorously validate these hypotheses, we conduct a comprehensive interpretability-driven analysis. Building upon these interpretability insights, we propose a minimal yet interpretable solution, namely, embedding norm regulation, which is aimed at explicitly restoring 2D spatial awareness in VLMs. Through controlled experiments, we demonstrate that embedding norm normalization effectively enhances spatial reasoning without compromising semantic understanding.
In summary, our contributions include:
  1. 1.
    An interpretability-driven theoretical explanation for VLMs’ spatial reasoning failures, rooted in embedding norm suppression.
  2. 2.
    Empirical interpretability evidence validating this theoretical insight through systematic permutation tests, embedding analysis, and attention visualization.
  3. 3.
    A simple, interpretable intervention that explicitly addresses spatial reasoning deficiencies, significantly improving performance on spatially demanding tasks.
The paper is organized as follows. After the introduction in Section 1, related work on VLMs, their intepretability and position encoding in Transformers is discussed in Section 2. Section 3 provides a diagnosis of the “Bag of Tokens” issue with token permutation and spatial compression. Detailed empirical and theoretical analyses are provided in Section 4 on spatial norm suppression. Section 5 proposes a controlled study in restoring spatial awareness of VLMs by designing a synthetic spatial benchmark. Experimental results on both synthetic and standard benchmarks are provided in Section 6. Section 7 analyzes how the combination of embedding normalization and multi-layer feature extraction influences the attention mechanisms of VLM. Section 8 discusses the findings as well as the limitations of the work, and Section 9 provides a few concluding remarks.
@@ 2 Related Work
Vision-Language Models Recent Vision-Language Models (VLMs) combine pretrained vision encoders and large language models (LLMs) to perform image understanding tasks, such as image captioning and visual question answering [23, 24, 20, 1, 41, 28, 7, 15]. Many approaches leverage pretrained visual encoders, notably CLIP [27], and subsequently introduce an adapter module to integrate vision embeddings into pretrained LLMs, such as Vicuna [6], enabling coherent textual outputs [23, 41]. While these methods achieve impressive performance in descriptive tasks, they often struggle with spatial reasoning [33, 38]. To address spatial limitations, recent efforts have focused on augmenting training datasets with extensive spatial annotations [32, 3, 5, 34]. In contrast, our approach does not rely on additional spatial labels or modifications to the original training data; instead, we introduce representation-level interventions, including embedding norm normalization and mid-layer feature extraction, which demonstrate significant improvements in VLMs’ spatial reasoning capabilities.
VLM Interpretability Interpretability of Vision-Language Models has attracted increasing attention, with the aim of explaining the internal mechanisms driving their successes or failures. Neo et al. [26] show that vision tokens in LLaVA [23] have strong semantic properties. Concurrently, multiple studies have highlighted the consistent failures of VLMs to comprehend simple spatial relations such as “above,” “below,” “left,” and “right” [38, 19, 33, 4]. These findings underscore a critical need for enhanced interpretability and robust spatial grounding in VLM architectures. Our work investigates the underlying reasons why spatial information is overshadowed in LLM, offering minimal yet effective interventions that notably strengthen the spatial reasoning ability.
Position Encoding in Transformers Positional encoding is essential in Transformer architectures to prevent degenerating into bag-of-words representations and to effectively model sequence orders [9, 35]. Initial efforts employed absolute positional encodings, later evolving into relative positional encodings [29] and, more recently, advanced mechanisms such as Rotary Positional Embeddings (RoPE) to better capture complex sequence dependencies [30]. Positional encoding has proven essential for modeling both language and visual data in Transformers [8]. Despite these developments, the specific impact of positional encodings within VLMs—where vision tokens interact with textual tokens in an LLM decoder—remains underexplored. Recent work [37] attributed hallucination issues to RoPE’s long-term decay property. Complementing these insights, our work empirically and theoretically illustrates how disproportionately large embedding norms of vision tokens can overshadow positional cues, hindering spatial reasoning even when advanced positional methods like RoPE are employed.
3 Diagnosing the Issue: Bag of Tokens
@@
VLMs demonstrate excellent capabilities in recognizing objects within images (ventral tasks) yet consistently struggle with spatial reasoning (dorsal tasks). This disparity suggests that VLMs might neglect spatial cues in their internal representations. Open-source VLMs such as LLaVA [23] typically fuse vision embeddings with textual embeddings in transformer-based language models, relying critically on positional embeddings to encode token order. We hypothesize that in multimodal contexts, positional embeddings might become ineffective or suppressed, causing a loss of meaningful spatial structure. To investigate this, we perform two diagnostic interpretability experiments: (1) Token Permutation Test: If positional order significantly contributes to spatial reasoning, randomizing token order should substantially degrade model performance. We want to know how much the vision token order matters. (2) Spatial Compression Study: If local spatial information is crucial, drastically reducing the number of vision tokens should cause significant performance deterioration. We want to know how much VLM really uses local spatial information. Our experiments indicate that VLMs remain largely robust to these two perturbations, reinforcing the hypothesis of a “bag-of-tokens” behavior.
3.1 Token Permutation Test
In standard NLP transformers, positional embeddings are essential for maintaining token order. Similarly, if VLMs depend heavily on positional embeddings for spatial reasoning, randomizing vision token order should substantially impair spatial reasoning performance. We hypothesize that if token order carries meaningful spatial information, random shuffling should severely degrade performance.
Experiment Setup: We use LLaVA 1.5 7B from [23, 24] to conduct the test. We randomly permute vision token embeddings obtained from the vision encoder and projection layers before feeding them into the LLM, and the performance is then compared against the original unpermuted baseline.  Figure 1 illustrates the experiment pipeline. We evaluate performance on datasets including VQAv2[11], POPE [21], GQA [16], and CV-Bench[32].
Result: As shown in  Table 1, random token permutation leads only to minor performance drops (ranging from 0.2% to 2.74%). This surprisingly small decline reveals that VLMs exhibit significant order-invariance, suggesting a bag of tokens tendency of VLMs. This is summarized as the following “Intepretabality Insight”:
Interpretability Insight No. 1:
VLMs exhibit limited sensitivity to vision token order, reflecting a strong “bag-of-tokens” tendency.
@@ 3.1 Token Permutation Test
3.2 Spatial Compression Study
Recent work showed VLMs focus on object semantic tokens[26]. To further probe the use of spatial information by VLMs, we progressively compress the vision tokens, effectively removing fine-grained spatial details including detailed positional and geometric relationships among image patches. If spatial details are crucial for evaluation, performance should sharply decline as spatial resolution decreases. Conversely, stable performance despite extreme compression may indicate either that the model relies predominantly on semantic features or that existing benchmarks do not effectively challenge spatial reasoning capabilities.
Experiment Setup: Using the LLaVA 1.5 7B model, we systematically reduce the number of vision tokens from the original 576 down to just 1. Models were retrained for token counts (256, 64, 16, and 1) using average pooling after the MLP projection layer. This method severely limits local spatial details while preserving some semantic content. Performance impacts are evaluated across various benchmarks.
Result: As illustrated in  Figure 2, reducing the vision tokens from 576 down to just 1 for standard benchmarks results in a minor accuracy drop (-8.5% the worst case). Despite losing approximately 99.8% of spatial resolution, the model’s performance remains robust. This outcome suggests two interrelated explanations: first, VLMs may primarily rely on semantic cues rather than local spatial details when solving typical benchmark tasks; second, the current benchmarks themselves might not effectively test the model’s capacity for detailed spatial reasoning. The limited sensitivity of these benchmarks to extreme spatial compression could reflect both model behavior and benchmark limitations. This leads us to the following insight:
Interpretability Insight No. 2:
For current benchmarks, VLMs can predominantly rely on “ventral” semantics rather than “dorsal” spatial cues to have a decent performance.
4 Why Does This Happen? Analysis of Embedding Norm Suppression
In earlier experiments, we observed that spatial information in VLMs tends to be minimal or overshadowed by strong global semantic signals. Here, we further investigate the underlying cause by analyzing embedding norms, specifically examining how large vision embedding magnitudes overshadow positional encodings within attention mechanisms.
4.1 Embedding Norm Analysis
Motivation Positional embeddings were initially introduced into Transformer architectures to maintain token order and avoid the “bag-of-words” effect [35]. While VLMs also utilize positional encodings during training, we hypothesize these encodings become ineffective if vision embeddings possess disproportionately large magnitudes, thus dominating positional signals during attention computations.
To investigate this hypothesis, we conduct an empirical embedding analysis using the COCO validation dataset (5000 image-text pairs) [22]. Specifically, we extract vision and text embeddings after MLP projection but before input into the language model and compute the L2 norms for each token type. Figure 3 illustrates the embedding norm distributions.
@@ 4.1 Embedding Norm Analysis
Figure 3: Distribution of L2 norms for vision and text tokens in COCO validation dataset (log scale). Vision token norms range between $10^{1}$ and $10^{3}$, while text token norms range between $3\times 10^{-1}$ and $10^{0}$.
As shown in Figure 3, vision embeddings consistently have 1-2 orders of magnitudes greater than text embeddings but can reach up to 3 orders of magnitude greater at its upper range. Although directional alignment (e.g., cosine similarity) typically receives primary focus in multimodal alignment due to inner product [14, 17, 25], such large magnitude discrepancies may suppress positional signals significantly.
4.2 Theoretical Explanation
We provide a concise theoretical justification of how differences in embedding norms significantly impact the effectiveness of positional encodings in Transformer attention mechanisms. A detailed mathematical derivation is presented in the Appendix A but here is a summary.
Summary of Proof (detailed derivation in Appendix A). Consider a Transformer self-attention block employing Rotary Positional Embeddings (RoPE) [31]. RoPE applies orthonormal rotations to query and key vectors with dimension $d$ ($\mathbf{q}_{i},\mathbf{k}_{i}\in\mathbb{R}^{d}$), preserving their norms. After RoPE transformation, we denote these vectors as $\mathbf{q}_{i}^{\prime},\mathbf{k}_{i}^{\prime}\in\mathbb{R}^{d}$. The attention logits between token pairs are computed via the scaled dot-product:
$\mathrm{logits}_{ij}=\frac{\mathbf{q}_{i}^{\prime}\cdot\mathbf{k}_{j}^{\prime}}{\sqrt{d}}=\frac{\|\mathbf{q}_{i}\|\|\mathbf{k}_{j}\|\cos{\phi}}{\sqrt{d}}$ (1)
where $\phi$ denotes the positional angle difference. From this expression, we observe directly that larger embedding norms increase the scale of logits. When vision embedding norms substantially exceed text embedding norms (often by an order of magnitude), vision-to-text logits dominate softmax computations, resulting in disproportionately high attention scores. To maintain coherent textual outputs aligned with the language model’s training, the model implicitly learns to downscale these vision-related logits. This downscaling leads to empirically observed low attention weights for vision tokens, consistent with findings in recent studies [4]. Consequently, the positional sensitivity explicitly encoded by RoPE diminishes substantially, as lower attention weights inherently reduce the positional gradient (further detailed in Appendix A).
Interpretability Insight No. 3 The practical effectiveness of positional encoding mechanisms’ (such as RoPE) is severely limited by embedding magnitude discrepancies between modalities.
Normalizing or rescaling embedding norms thus emerges as a crucial intervention, restoring positional sensitivity and enabling positional embeddings to meaningfully influence attention computations. This insight extends broadly to multi-modal alignment scenarios: effective alignment between pretrained models should not rely solely on embedding similarities (such as cosine similarity) but must also carefully consider embedding norm differences. Otherwise, positional information can become suppressed, leading to minimal influence on model predictions.
5 Restoring Spatial Awareness: A Controlled Study
Our previous analyses indicated that spatial cues in VLMs are often overshadowed by strong semantic signals and disproportionately large embedding norms. To validate these findings and explore methods for enhancing spatial representations, we designed a controlled study employing two interpretability-informed strategies:
  1. 1.
    Synthetic Dataset: A carefully constructed two-dimensional (2D) spatial benchmark designed explicitly to eliminate semantic shortcuts, ensuring that spatial relationships alone drive task success.
  2. 2.
    Interpretability-Informed Model Adjustments: Targeted architectural modifications aimed at balancing positional cues and mitigating the overwhelming dominance of semantic embeddings.
Combining these controlled experimental conditions with interpretability-driven adjustments allows us to investigate the feasibility of reviving effective spatial reasoning—analogous to restoring a functional “dorsal stream”—in VLMs.

### v2

@@ Abstract
Vision–Language Models (VLMs) excel at identifying and describing objects but often fail at spatial reasoning. We study why VLMs, such as LLaVA, under-utilize spatial cues despite having positional encodings and spatially rich vision-encoder features. Our analysis reveals a key imbalance: vision token embeddings have much larger norms than text tokens, suppressing LLM’s position embedding. To expose this mechanism, we developed three interpretability tools: (1) the Position Sensitivity Index, which quantifies reliance on token order, (2) the Cross-Modality Balance, which reveals attention head allocation patterns, and (3) a RoPE Sensitivity probe, which measures dependence on rotary positional embeddings. These tools uncover that vision tokens and system prompts dominate attention. We validated our mechanistic understanding through targeted interventions that predictably restore positional sensitivity. These findings reveal previously unknown failure modes in multimodal attention and demonstrate how interpretability analysis can guide principled improvements. Website is at: user074.github.io/respatialaware/.
Figure 1: Permutation Test: Original (top) vs. randomly permuted vision tokens (bottom). Despite losing spatial ordering, the LLM accurately responds to the prompt “Describe the image,” demonstrating strong robustness and a notable “bag-of-tokens” tendency, losing spatial relationships. Token embeddings are visualized using cosine similarity relative to a reference token.
@@ 1 Introduction
Humans process vision through two distinct pathways: the ventral stream, which identifies objects, and the dorsal stream, which encodes spatial relationships. The dorsal stream, located in the parietal lobe, is essential for processing spatial relationships, allowing us to understand where objects are and how they relate to each other. Not surprisingly, patients with damage to the parietal lobe exhibit asymmetric visual processing: they retain the ability to recognize and describe objects but lose spatial reasoning capabilities (Goodale & Milner, 1992).
@@
Recent observations suggest that generative Vision-Language Models (VLMs) exhibit a similar asymmetry. Despite remarkable performance in object-centric (ventral) tasks such as object recognition and image captioning (Goyal et al., 2017; Antol et al., 2015; Gurari et al., 2018), VLMs struggle with even fundamental spatial reasoning queries (e.g., distinguishing “left” from “right”) (Kamath et al., 2023; Tong et al., 2024b). This leads to a critical interpretability question: Where does “space” live in VLMs, and why is it under-used?
In this paper, we only focus on the common architecture that couples a pre-trained vision encoder with a pre-trained language model (LLM). In principle, both sides provide spatial signals: vision encoders carry layout through their feature maps and token order, while the LLM’s positional mechanism (RoPE) can make the decoder sensitive to the token order. Yet the combined system often behaves like a “bag of words”, mostly losing spatial relationships (Yuksekgonul et al., 2022). So the question is: Why does the integration fail?
We hypothesize two causes: First, the RoPE is under-utilized. Second, the last few layers of the vision encoder might not contain enough spatial information. We first diagnose and show empirical evidence of the norm mismatch of the two modalities (vision and text) causing attention imbalance. Then we develop a compact interpretability toolkit and two mechanism-aligned interventions. Our toolkit comprises: (i) a Position Sensitivity Index (PSI) tool that quantifies order use, (ii) a Cross-Modality Balance (CMB) tool to analyze head-level vision vs. text attention with prompt isolation, and (iii) a RoPE sensitivity probe tool that measures intrinsic response to positional phase shifts. We also introduce a controlled synthetic benchmark (2DS) to better measure these three indicators.
We then apply two minimal interventions to test our hypothesis: Normalize RMS scale-matching of vision embeddings post-projector, and Normalize+Multilayer combination of intermediate encoder layers. We do an extensive comparison with our interpretability tools to understand how our interventions can make the model behave differently. Our aim is understanding and measurement, instead of advancing SOTA.
Contributions. (1) Diagnosis: we identify and validate a concrete mechanism —embedding-norm skew— that links modality scale to diminished positional sensitivity in the LLM. (2) Toolkit & data: PSI, CMB, and RoPE probe, together with a spatially focused dataset 2DS offer complementary views of spatial use. (3) Mechanism-aligned probes: we demonstrate that norm-induced problem is the causal mechanism behind positional insensitivity and that intermediate layers of the vision encoder can contribute more to spatial understanding.
@@ 2 Related Work
Vision-Language Models Recent Vision-Language Models (VLMs) combine pretrained vision encoders and large language models (LLMs) to perform image understanding tasks, such as image captioning and visual question answering (Liu et al., 2023; 2024; Li et al., 2023a; Alayrac et al., 2022; Zhu et al., 2024; Shao et al., 2023; Ding et al., 2022; Hu et al., 2022). Many approaches leverage pretrained visual encoders, notably CLIP (Radford et al., 2021), and subsequently introduce an adapter module to integrate vision embeddings into pretrained LLMs, such as Vicuna (Chiang et al., 2023), enabling coherent textual outputs (Liu et al., 2023; Zhu et al., 2024). While these methods achieve impressive performance in descriptive tasks, they often struggle with spatial reasoning (Tong et al., 2024b; Yuksekgonul et al., 2023). To address spatial limitations, recent efforts have focused on augmenting training datasets with extensive spatial annotations (Tong et al., 2024a; Chen et al., 2024a; Cheng et al., 2024; Trabelsi et al., 2025). In contrast, our approach does not rely on additional spatial labels or modifications to the original training data; instead, we introduce representation-level interventions, including embedding norm normalization and mid-layer feature extraction, which demonstrate significant improvements in VLMs’ spatial reasoning capabilities.
VLM Interpretability Interpretability of Vision-Language Models has attracted increasing attention, with the aim of explaining the internal mechanisms driving their successes or failures. Neo et al. (2025) show that vision tokens in LLaVA (Liu et al., 2023) have strong semantic properties. Concurrently, multiple studies have highlighted the consistent failures of VLMs to comprehend simple spatial relations such as “above,” “below,” “left,” and “right” (Yuksekgonul et al., 2023; Kamath et al., 2023; Tong et al., 2024b; Chen et al., 2025). These findings underscore a critical need for enhanced interpretability and robust spatial grounding in VLM architectures. Our work investigates the underlying reasons why spatial information is overshadowed in LLMs, offering minimal yet effective interventions that notably strengthen the spatial reasoning ability.
3 Motivation
@@
VLMs demonstrate excellent capabilities in recognizing objects within images (ventral tasks) yet consistently struggle with spatial reasoning (dorsal tasks). Open-source VLMs such as LLaVA (Liu et al., 2023) typically consist of a vision encoder and an LLM. One central question we want to know is: How do VLMs use the spatial information? Conceptually, there are two sources for the spatial information: vision encoder and LLM position embedding. The vision encoder encodes spatial information within vision tokens. LLM position embedding directly represents the spatial information of the input tokens’ order. We are particularly interested in how much the LLM leverages position embedding for spatial information.
To understand the positional sensitivity, we use LLaVA 1.5 7B (Liu et al., 2023; 2024) to randomly permute vision token embeddings after processing by the vision encoder and the MLP projection layers, but right before they are fed into the LLM (Figure 1). It effectively changes the position embedding for each token. The performance is then compared against the original unpermuted baseline. We evaluate performance on datasets including VQAv2 (Goyal et al., 2017), POPE (Li et al., 2023c), GQA (Hudson & Manning, 2019), and CV-Bench (Tong et al., 2024a). As shown in Table  1, random token permutation leads only to minor performance drops (ranging from 0.2% to 2.74%). This surprisingly small decline reveals that current VLMs evaluated with the corresponding benchmarks are largely insensitive to positional information.
@@
Another way to investigate the role of position is to eliminate the position embedding completely. Recent work showed VLMs focus on object semantic tokens (Neo et al., 2025), and we want to know how much spatial information is encoded into the vision tokens and how much is the result of the LLM position embedding. We progressively compress the vision tokens, effectively removing the position embedding’s effect on the spatial information, including detailed positional and geometric relationships among image patches. We aim to push it to the extreme to see how much information is relied on the LLM position embedding or if they are already encoded in the vision token. Using the LLaVA 1.5 7B model, we retrained with vision token counts of 256, 64, 16, and 1 (originally 576), using average pooling after the MLP projection layer. As illustrated in Figure  2, reducing the vision tokens from 576 down to just 1 for standard benchmarks results in a minor accuracy drop (-8.5% the worst case). Despite losing approximately 99.8% (from 576 to 1) of spatial resolution, the model’s performance remains robust. Concurrent work also reached similar conclusion and showed robustness of compression (Chai et al., 2025).
This outcome suggests two possible reasons. First, VLMs may primarily rely on vision encoder’s spatial encoding, and position embedding in the LLM has low effects. Second, the current benchmarks might not effectively test the model’s capacity with the LLM part of the spatial information encoding.
4 Why Does This Happen? Embedding Norm Suppression
The previous section showed a striking failure of positional encoding: permuting vision tokens often leaves predictions minor change, i.e., an unintended “bag-of-words” behavior despite the presence of positional encodings (Vaswani et al., 2017; Yuksekgonul et al., 2022). To probe why, we conduct an empirical norm analysis on COCO validation (5k image–text pairs) (Lin et al., 2015). We extract vision embeddings after the MLP projector and before the language model, compute their $L_{2}$ norms, and compare to text-token norms. Figure 3 shows the distribution comparison.
@@ 4 Why Does This Happen? Embedding Norm Suppression
Figure 3: Distribution of $L_{2}$ norms for vision and text tokens in COCO validation dataset (log scale). Vision token norms range between $10^{1}$ and $10^{3}$, while text token norms range between $3\times 10^{-1}$ and $10^{0}$.
In LLaVA 1.5 model, vision embeddings are typically 1–2 orders of magnitude larger than text embeddings, with an upper tail approaching 3 orders. While alignment work often emphasizes direction (cosine similarity) because logits are inner products (Hessel et al., 2021; Huh et al., 2024; Maniparambil et al., 2024), such large magnitude gaps can dominate the scaled dot-product and suppress positional signals.
Our analysis focuses on the LLaVA architecture, which uses a LLaMA decoder (Grattafiori et al., 2024). LLaMA employs a pre-Norm design where each sublayer follows a specific pattern: (1) the input hidden states is first normalized by RMSNorm (Zhang & Sennrich, 2019b), (2) the computational block (e.g., self-attention) processes this normalized input to produce an update, and (3) this update is added back to the original, unnormalized residual. RoPE is applied within the self-attention block to the normalized inputs (Su et al., 2024). We hypothesize that the large vision norms in the residual path are not normalized. It reduces the RoPE effectiveness applied in the attention when we consider the magnitude difference between attention outputs and unnormalized vision tokens. It might also let the vision token direction dominate due to small changes to the residual in early layers, which could cause vision attention bias.
[…]
If the above mentioned hypothesis is correct, then the hidden states of vision tokens should be much larger than text tokens in early layers. On COCO val (5k), we partition tokens into vision ($V$) and user text ($T$). We exclude the fixed system prompt for clarity. We compute each token’s $L_{2}$ norms and average over respective tokens. Figure 4 shows the results. Vision hidden states are an order of magnitude larger than text in early layers, and the disparity persists until roughly mid‑depth (Figure 4 right). This directly supports the residual‑scale suppression account.
More detailed explanation is given in Appendix A. Overall we want to measure the RoPE effectiveness directly to see its impact. To measure the RoPE effect directly, we can check the partial derivative of the attention w.r.t. position phase ($\phi$)’s changes. Let logits be $\ell_{j}=\langle q^{\prime},k^{\prime}_{j}\rangle/\sqrt{d}$ and $\alpha_{k}=\mathrm{softmax}(\ell)_{k}$ be attention weights, and partition tokens into vision ($V$) and text ($T$). Define the group masses $\alpha_{V}=\sum_{v\in V}\alpha_{v}$, $\alpha_{T}=\sum_{t\in T}\alpha_{t}$, and the group‑average logit derivatives
$g_{V}\;=\;\sum_{v\in V}\frac{\alpha_{v}}{\alpha_{V}}\,\frac{\partial\ell_{v}}{\partial\phi},\qquad g_{T}\;=\;\sum_{t\in T}\frac{\alpha_{t}}{\alpha_{T}}\,\frac{\partial\ell_{t}}{\partial\phi}.$
Then (derivation in Appendix A) the change of total vision attention mass w.r.t. the positional phase $\phi$ is
$\frac{\partial\alpha_{V}}{\partial\phi}\;=\;\alpha_{V}\,\alpha_{T}\,\bigl(g_{V}-g_{T}\bigr).$ (1)
This derivation establishes a direct link between the shift in vision attention mass ($\frac{\partial\alpha_{V}}{\partial\phi}$) and the underlying positional sensitivity of the vision logits ($g_{V}$). According to this relationship, a very small attention shift would mean low positional sensitivity. Conversely, a large attention shift would indicate high positional sensitivity, suggesting that RoPE is functioning effectively. This principle provides the mathematical support for RoPE sensitivity probe in the next section.
5 Interpretability Toolkit: Understand who is looking at what
We develop a compact toolkit to probe where spatial information is (or is not) used inside VLMs. The tools target three complementary questions: (i) Does position embedding matter? (Position Sensitivity), (ii) Which components actually look at vision? (Cross-Modality Balance at the head level), and (iii) How sensitive is the model to positional phase shifts? (RoPE Sensitivity).
5.1 Position Sensitivity Index (PSI)
As discussed in Section 3, shuffling vision tokens often leaves answers mostly unchanged, suggesting limited use of spatial order through position embedding. We quantify this with the Position Sensitivity Index:
$\mathrm{PSI}\;=\;\frac{\mathrm{Acc}(\text{original order})-\mathrm{Acc}(\text{permuted order})}{\mathrm{Acc}(\text{original order})}.$ (2)
PSI is simple, scale-free, and interpretable: $\mathrm{PSI}=0$ means complete order invariance, and larger values indicate greater reliance on the token order. We report PSI at both the model and dataset levels since it could be both sides’ issues. This lets us compare models and datasets on a common measure.
5.2 Cross Modality Balance (CMB)
Functional behavior in Transformers is often head-specific, for example there are induction, copy, memory, summary, truthfulness attention heads in LLM (Zheng et al., 2024; Bietti et al., 2023; Tigges et al., 2023; Olsson et al., 2022; Li et al., 2023b). We therefore work at head granularity to give a closer look at the vision and text modality balance.
Let $V$ and $T$ denote the index sets of vision and text tokens, respectively. For a fixed decoding step and attention head $h$ with attention weights $a_{\cdot,h}$, we have:
$\mathrm{CMB}_{h}\;=\;\frac{\sum_{v\in V}a_{v,h}}{\sum_{i\in V\cup T}a_{i,h}}.$ (3)
where $\mathrm{CMB}_{h}\in[0,1]$ measures how much attention head $h$ attends to vision versus text. In practice, we summarize the CMB with layer-by-layer head heatmaps. We can visualize which head is more focused on the vision and which is more focused on the text.
5.3 RoPE Sensitivity Probe
To empirically quantify the practical impact of RoPE’s positional information on the models’ attention mechanism, we introduce a RoPE sensitivity probe to measure how much positional phase (RoPE) affects what the model predicts. For a fixed query position, we first extract the query vector and all key vectors to compute the baseline attention score, denoted as $\alpha$. We then create a perturbed state by applying an additional RoPE rotation of $\Delta$ steps exclusively to vision keys. The query vector and text key vectors remain unchanged and preserve original causal masking. Finally, we recompute the attention score, $\alpha_{\Delta}$, using the perturbed vision keys. This perturbation allows us to isolate how the positions of visual evidence shift relative to a fixed linguistic query. From this procedure, we report two key measures that serve as a finite-difference approximation of terms in Equation 4, which is also listed here for its importance:
$\frac{\partial\alpha_{V}}{\partial\phi}\;=\;\alpha_{V}\,\alpha_{T}\,\bigl(g_{V}-g_{T}\bigr).$ (4)
(A) Attention-level sensitivity. As an approximation of $\frac{\partial\alpha_{V}}{\partial\phi}$, $\Delta\alpha_{V}$ is defined as the finite difference between $\alpha$ and $\alpha_{\Delta}$ with step $\Delta$. This captures the observed change in how much attention mass moves among vision tokens under a RoPE phase shift.
(B) Intrinsic (logit-level) sensitivity. We can also directly measure $\Delta g_{V}$ which is the group-average logit derivative w.r.t. the RoPE angle. It is an approximation of the $g_{V}$ term. In experiments we report both the attention-level difference and the normalized logit-level difference.